A plate glass window (n = 1.5) has a thickness of 3.7 mm. How long does it take light to pass perpendicularly through the plate?
Answer #1

Velocity of light in glass = c/n

= 3*10^8 / 1.5

= 2*10^8

Time = thickness/ speed

= 3.7*10^-3 / 2*10^8

= 1.85 *10^-11 sec